OpenStudy (anonymous):

Which is a solution of the equation? 2x – y = 5 A. 7 B. (1, 2) C. (3, –1) D. (2, –1)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I think it is D

OpenStudy (anonymous):

2(2)-y=5 4-y=5 -y=5-4 y=-1

OpenStudy (anonymous):

(3,-1) is different from (1,-3)